OpenStudy (anonymous):

What polynomial has roots of -4, 1, and 6 ? x3 - 3x2 - 22x + 24 x3 - x2 - 26x - 24 x3 + x2 - 26x + 24 x3 + 3x2 + 14x - 24 i need an answer quick i dont have much time :/

OpenStudy (anonymous):

(x + 4)(x - 1)(x - 6) = 0 (x + 4)(x^2 - 7x + 6) = 0 x^3 - 7x^2 + 6x + 4x^2 - 28x + 24 = 0 x^3 - 3x^2 - 22x + 24 = 0